tTwo astronauts, each having a mass of 38.0 kg, are connected by a 10.0-m rope of regligible maes. They are sclated in space,two astronauts, each having a mass of 88.0 kg, are connected by a 10.0-m rope of negligible mass. They are isolated in space, moving in circles around the point halfway between them at a speed of 5.60 m/s. Treating the astronauts as particles, calculate each of the following.

Two astronauts are connected by a taut horizontal rope of length d. They rotate counterclockwise about a point labeled CG at the midpoint of the rope.

(a) the magnitude of the angular momentum of the system
kg · m2/s

(b) the rotational energy of the system
kJ


By pulling on the rope, the astronauts shorten the distance between them to 5.00 m.

(c) What is the new angular momentum of the system?
kg · m2/s

(d) What are their new speeds?
m/s

(e) What is the new rotational energy of the system?
kJ

(f) How much work is done by the astronauts in shortening the rope?
kJ
Answer #1

m=88kg ,dziom u=5.6 m/s ,r=d2=5m a) L;= mortmur li= 2mur = 2 *88x5.6*5 = 4928 kg mºls b) K; = 1 I w² + + Iw² = 2x(mx r²) x 2